# A Cold and Super-Puffy Planet on a Polar Orbit

### v1

@@ 5 Joint Photometric and RV Modeling
$\psi$ True 3D stellar obliquity (deg) $82.0_{-2.4}^{+2.6}$
$\lambda$ Sky-projected stellar obliquity (deg) $\mathcal{U}(-180,180)$ $49_{-28}^{+51}$
$v$ Equatorial velocity (km/s) $29.4_{-2.2}^{+2.3}$
$v\sin{i_{\star}}$ Projected rotational velocity (km/s) $7.1_{-2.0}^{+6.1}$
$\cos{i_{\star}}$ Cosine of stellar inclination $\mathcal{U}(0,1)$ $0.97_{-0.08}^{+0.02}$
$i_{\star}$ Stellar inclination (deg) $14.1_{-4.2}^{+12.7}$
$P_{\rm rot}$ Stellar rotational period (days) $\mathcal{G}(1.70,0.17)$ $1.79_{-0.13}^{+0.14}$
$M_{\star}$ Stellar mass ($M_{\odot}$) $\mathcal{G}(1.11,0.02)$ $1.11\pm 0.02$
$R_{\star}$ Stellar radius ($R_{\odot}$) $\mathcal{G}(1.04,0.01)$ $1.04\pm 0.01$
$\rho_{\star}$ Stellar density (g/cm3) $1.37\pm 0.04$
$P$ Orbital period (days) $\mathcal{G}(104.61573,0.00008)$ $104.6159518\pm 0.00004$
$t_{0}$ Transit midpoint (BJD) $\mathcal{G}(2458413.8542,0.0007)$ $2458413.85343\pm 0.0005$
$b$ Impact parameter $\mathcal{U}(0,1)$ $0.08_{-0.06}^{+0.08}$
$i$ Orbital inclination (deg) $89.95_{-0.05}^{+0.04}$
@@
$a/R_{\star}$ Scaled semimajor axis $93\pm 1$
$K$ RV semiamplitude (m/s) $\mathcal{U}(-10,10)$ <4<4 ($2\sigma$)
$\sqrt{e}\sin{\omega}$ Eccentricity parameter — sine component $\mathcal{U}(-1,1)$ $0.09_{-0.05}^{+0.06}$
$\sqrt{e}\cos{\omega}$ Eccentricity parameter — cosine component $\mathcal{U}(-1,1)$ $0.11_{-0.29}^{+0.25}$
$e$ Eccentricity $0.05_{-0.04}^{+0.12}$
$\omega$ Argument of periastron (deg) $32_{-125}^{+23}$
$M_{p}$ Planet mass ($M_{\oplus}$) <30<30 ($2\sigma$)
$R_{p}$ Planet radius ($R_{\oplus}$) $8.22\pm 0.08$
$\rho_{p}$ Planet mean density (g/cm3) <0.3<0.3 ($2\sigma$)
$a$ Semimajor axis (au) $0.449\pm 0.003$
@@
$\ddot{\gamma}$ RV quadratic trend (m/s/day/yr) $\mathcal{U}(-1,1)$ $0.0183\pm 0.0036$
$\gamma_{\rm HARPS1}$ HARPS1 RV offset (m/s) $\mathcal{U}(-100,100)$ $20\pm 4$
$\sigma_{\rm HARPS1}$ HARPS1 RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $11.8_{-2.4}^{+3.2}$
$q_{1}^{\rm HARPS2}$ HARPS2 linear limb darkening parameter $\mathcal{U}(0,1)$ $0.51_{-0.29}^{+0.30}$
$q_{2}^{\rm HARPS2}$ HARPS2 quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.61_{-0.34}^{+0.27}$
$\beta_{\rm HARPS2}$ Intrinsic stellar line width (km/s) $\mathcal{G}(6.1,2.0)$ $6.8_{-1.8}^{+1.7}$
$\gamma_{\rm HARPS2}$ HARPS2 RV offset (m/s) $\mathcal{U}(-100,100)$ $55\pm 10$
$\sigma_{\rm HARPS2}$ HARPS2 RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $9.0_{-1.1}^{+1.3}$
$\gamma_{\rm FEROS}$ FEROS RV offset (m/s) $\mathcal{U}(25000,27000)$ $26062_{-9}^{+8}$
$\sigma_{\rm FEROS}$ FEROS RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $19.9_{-6.0}^{+7.8}$
$\gamma_{\rm CORALIE}$ CORALIE RV offset (m/s) $\mathcal{U}(25000,27000)$ $26076\pm 9$
$\sigma_{\rm CORALIE}$ CORALIE RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $0.1_{-0.1}^{+2.6}$
$q_{1}^{\rm TESS}$ TESS linear limb darkening parameter $\mathcal{U}(0,1)$ $0.42_{-0.11}^{+0.12}$
$q_{2}^{\rm TESS}$ TESS quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.11_{-0.07}^{+0.10}$
$\sigma_{\rm TESS}$ TESS photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $6_{-4}^{+17}$
$q_{1}^{R}$ $R$ linear limb darkening parameter $\mathcal{U}(0,1)$ $0.52_{-0.06}^{+0.08}$
$q_{2}^{R}$ $R$ quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.92_{-0.10}^{+0.06}$
$\sigma_{\rm ASTEP}^{2023-R}$ ASTEP 2023 $R$ photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $3098_{-75}^{+79}$
$\sigma_{\rm ASTEP}^{2025-R}$ ASTEP 2025 $R$ photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $8969_{-88}^{+84}$
$q_{1}^{B}$ $B$ linear limb darkening parameter $\mathcal{U}(0,1)$ $0.17_{-0.12}^{+0.20}$
$q_{2}^{B}$ $B$ quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.21_{-0.16}^{+0.33}$
$\sigma_{\rm ASTEP}^{2023-B}$ ASTEP 2023 $B$ photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $2250_{-92}^{+103}$
Note. — $\mathcal{U}(a,b)$ denotes a uniform prior with a start value $a$ and end value $b$. $\mathcal{G}(\mu,\sigma)$ denotes a Gaussian prior with mean $\mu$, and standard deviation $\sigma$. $\mathcal{LU}(a,b)$ denotes a log-uniform prior with a start value $a$ and end value $b$.
@@ 6 Discussion
6.1 Is this a truly polar orbit?
While we detected the RM effect, which validates the candidate with photometric and spectroscopic detections of its transits (e.g., B. S. Gaudi & J. N. Winn, 2007), we acknowledge that the result of TOI-4507 b having a polar orbit may be questionable. We did not observe a full transit, so the measured sky-projected obliquity $\lambda=-49_{-28}^{+51}$ deg has a large error bar that is only $\sim 1\sigma$ away from an aligned orbit and $\sim 1.5\sigma$ away from polar. The conclusion of a polar orbit comes from the $v\sin{i_{\star}}$ and stellar rotational period values, and the reliability of these measurements.
The stellar analysis presented in Section 3, which estimates $v\sin{i_{\star}}$ using the broadening of the spectral lines, resulted in $v\sin{i_{\star}}=4.6\pm 0.9$ km/s. Furthermore, Gaia DR3 reports a spectral line broadening of $5.2\pm 4.4$ km/s. Consistently, the RM analysis, which estimates $v\sin{i_{\star}}$ using the amplitude of the RM effect, resulted in $v\sin{i_{\star}}=7.1_{-2.0}^{+6.1}$ km/s. We argue that the $v\sin{i_{\star}}$ measurement is then solid, as the values derived with independent methodologies are in good agreement.
As for the rotational period, we attempted to validate it using long-term ground-based photometric monitoring. Unfortunately, the star was not observed by the Zwicky Transient Facility (E. C. Bellm et al., 2019), and our analysis of the ASAS-SN (B. J. Shappee et al., 2014; C. S. Kochanek et al., 2017) light curve only recovered the well-known 1-day alias typically seen in ground-based data (e.g., J. T. VanderPlas, 2018). The confidence in the rotational period measurement is then only based on TESS data. A. W. Boyle et al. (2025) recently studied the limits of TESS in measuring rotational periods using the overlap between TESS and K2. They concluded that rotational periods shorter than 5 days derived from a simple Lomb-Scargle of TESS data are good to $\sim 1-2\%$, with minimal $\sim 10\%$ alias detection. Therefore, our detection of a periodic signal with a period of 1.7 days is solid.
If not related to the stellar rotation, the observed periodic signal might be caused by pulsations. The short period of only 1.7 days and the F-type classification of the host star suggest that it could be a $\gamma$ Doradus variable (e.g., C. Aerts et al., 2004). Indeed, the photometric variability amplitude of $\sim 500$ ppm lies within the expected range for this class of pulsators, which typically range from a few hundred to tens of thousands of ppm (e.g., L. A. Balona et al., 2011). However, the star is relatively cool to be a bona fide $\gamma$ Doradus variable, and it would lie outside their typical empirical instability strip (e.g., L. A. Balona, 2014; F. Kahraman Aliçavuş et al., 2016). Additionally, the variability amplitude appears relatively constant across cycles, in contrast to the cycle-to-cycle amplitude modulations commonly observed in $\gamma$ Doradus stars (e.g., V. Antoci et al., 2019; M. Skarka et al., 2022; M. Skarka & Z. Henzl, 2024). Furthermore, at least some $\gamma$ Doradus are known to produce RV variations on the order of a few km/s per cycle (e.g., G. Handler et al., 2002; C. Aerts et al., 2004; G. W. Henry et al., 2022), whereas TOI-4507 does not show any variation larger than $\sim 10$ m/s. The pulsation hypothesis is thus disfavored, supporting the interpretation that the observed variability is due to rotation. Consequently, this strengthens our conclusion that TOI-4507 b is on a nearly polar orbit.
@@ 6.2 TOI-4507 b in context
TOI-4507 b appears to be an interesting planet in several ways. Figure 4 shows the results for the obliquity $\psi$ and its sky projection $\lambda$ along with those of other systems. The obliquity data were drawn from TEPCat [^3] https://www.astro.keele.ac.uk/jkt/tepcat/ (J. Southworth, 2011) and the age data were drawn from the NASA Exoplanet Archive (R. L. Akeson et al., 2013; J. L. Christiansen et al., 2025), both as of May 2025. Evidently, TOI-4507 b has the longest orbital period of any planet for which $\psi$ has been measured, and has the third-longest orbital period of any planet for which $\lambda$ has been measured [after HIP 41378 d (S. Grouffal et al., 2022), TIC 241249530 b (A. F. Gupta et al., 2024), and HD 80606 b (e.g., G. Hébrard et al., 2010)]. Here, we define Neptunes/sub-Saturns as the class of planets with $10<M_{p}/M_{\oplus}<90$, with the subclass of super-puffs having $\rho_{p}<0.3$ g/cm3.
TOI-4507 b joins a growing population of Neptunes/sub-Saturns that are observed to have nearly polar orbits (e.g., J. I. Espinoza-Retamal et al., 2024; E. Knudstrup et al., 2024; L. B. Handley et al., 2025). This population includes some other super-puffs such as KELT-11 b (D. Mounzer et al., 2022), WASP-107 b (F. Dai & J. N. Winn, 2017; R. A. Rubenzahl et al., 2021; V. Bourrier et al., 2023), WASP-127 b (R. Allart et al., 2020; E. Cristo et al., 2022), and WASP-131 b (L. Doyle et al., 2023; J. Zak et al., 2024). Within this collection of polar super-puffs, TOI-4507 b has the longest period.
Figure 4: Sky-projected obliquity $\lambda$ (top panels) and 3D obliquity $\psi$ (bottom panels) as a function of the orbital period (left panels) and age of the system (right panels). The population of Neptunes/sub-Saturns ($10<M_{p}/M_{\oplus}<90$) is shown in blue. Super-puff planets (Neptunes/sub-Saturns with $\rho_{p}<0.3$ g/cm3) are shown in pink, with TOI-4507 b highlighted as a star with a red edge. Other systems are shown as faint black points. Data from TEPCat with ages from the NASA Exoplanet Archive.
Additionally, with an age of $\sim 700$ Myr, TOI-4507 b is the youngest super-puff with an obliquity measurement, and one of the youngest members of the polar Neptune/sub-Saturn population. The only polar Neptune younger than it is possibly AU Mic c (E. Martioli et al., 2021), for which the projected obliquity has been reported as $\lambda=68^{+32}_{-49}$ deg (H. Yu et al., 2025), but the result in that case is tentative and might not be compatible with long-term dynamical stability. More generally, TOI-4507 b is one of the youngest polar planets known of any size or density, comparable to Kepler-63 b (R. Sanchis-Ojeda et al., 2013; V. Bourrier et al., 2023), KELT-9 b (e.g., B. S. Gaudi et al., 2017; J. P. Ahlers et al., 2020), and KELT-17 b (G. Zhou et al., 2016).
@@
TOI-4507 b seems well-suited for atmospheric characterization via transit spectroscopy. Assuming zero albedo, the expected equilibrium temperature of the planet is $\sim 460$ K, lower than most other systems that have been subject to transit spectroscopy. Because of the planet’s low density, the transmission spectroscopy metric (TSM; E. M. R. Kempton et al., 2018) is $\gtrsim 100$. For reference, targets with TSM > 90 are typically considered high-quality targets for atmospheric characterization. Observations with JWST might shed light on the atmospheric composition of TOI-4507 b and the origins of its unusually low density.
@@ 6.3 Search for companions
If the RV trend is produced by a planetary companion, it should have an orbital period $P>5.6$ yr, which corresponds to the baseline of the current RV observations. The parameters of this possible companion could be constrained with additional RV and/or Gaia DR4 astrometric measurements. When removing the signal of TOI-4507 b and the RV trend from the data, we do not see clear additional signals in the RV data, which rules out the presence of planets producing similar or larger RV signals than TOI-4507 b.
@@
6.4 Orbital architecture
Here, we discuss some of the proposed mechanisms to explain polar orbits and whether they can explain the observed properties of the TOI-4507 system.
Primordial disk misalignment
The highly misaligned and almost circular orbit of TOI-4507 b can be naturally explained by a primordial misalignment between the star and the protoplanetary disk where the planet formed. However, most of the proposed mechanisms to tilt disks invoke gravitational torques from binary companions (e.g., K. Batygin, 2012; D. Lai, 2014; J. J. Zanazzi & D. Lai, 2018), while we found no evidence of such stellar companions. An alternative mechanism could be stellar fly-bys (e.g., N. Cuello et al., 2019) or a magnetic torque between the inner disk and the star, which could have inclined the star with respect to the disk (e.g., D. Lai et al., 2011; C. Spalding & K. Batygin, 2015).
Planet-planet scattering
The misaligned orbit of TOI-4507 b could have resulted from planet-planet scattering near its current orbital separation. In principle, scattering can produce large misalignments by ejecting planets. However, such events also tend to excite high eccentricities, in contrast with the observed small value. Moreover, the ratio between the escape velocity of the planet and its orbital velocity at the current semimajor axis is $\sim 0.1$, implying that scattering in this regime leads primarily to collisions rather than ejections (C. Petrovich et al., 2014). Such collisional outcomes typically leave the surviving planet on a low-eccentricity and low-obliquity orbit, making this scenario unlikely to explain the observed nearly polar orbit.
Excitation by a distant companion
Another possibility is that an outer, inclined companion secularly excites the obliquity through nodal precession of the orbital plane. Explaining the nearly polar orbit of TOI-4507 b requires a mutual inclination between the planetary orbits $\gtrsim 40$ deg (because $\psi$ oscillates between 0 and twice the mutual inclination). In this regime, TOI-4507 b may undergo von Zeipel-Lidov-Kozai oscillations (H. von Zeipel, 1910; M. L. Lidov, 1962; Y. Kozai, 1962), but this would strongly excite its orbital eccentricity, making it difficult to reconcile with the low observed eccentricity, unless there is another companion that suppresses eccentricity growth.
Alternatively, a planetary companion can also secularly excite large stellar obliquities by an inclination secular resonance during the disk dispersal phase (C. Petrovich et al., 2020). However, nearly polar orbits are only possible when relativistic precession dominates over the precession induced by the planetary companion, which is unlikely given the long orbital period of the planet [^4] Assuming an outer planetary companion with mass $3M_{J}$, the precession rates become comparable if its semi-major axis is $\gtrsim 13$ au. The secular timescale due to such a distant companion may be likely long ($\gtrsim 0.4$ Myr) to allow for an adiabatic capture..

### v2

@@ 6 Rossiter-McLaughlin Effect
$\rho_{\star}$ Stellar density (g/cm3) $\mathcal{G}(1.40,0.05)$ $1.41\pm 0.04$
$P$ Orbital period (days) $\mathcal{U}(104.61493,104.61653)^{a}$ $104.61611\pm 0.00006$
$t_{0}$ Transit midpoint (BJD - 2,458,000) $\mathcal{U}(413.8472,413.8612)^{a}$ $2458413.8515\pm 0.0008$
$b$ Impact parameter $\mathcal{U}(0,1)$ $0.12\pm 0.08$
$i$ Orbital inclination (deg) $89.93\pm 0.05$
@@
$a/R_{\star}$ Scaled semimajor axis $93.4_{-0.9}^{+0.8}$
$K$ RV semiamplitude (m/s) $\mathcal{U}(-100,100)$ <3<3 ($2\sigma$)
$\sqrt{e}\sin{\omega}$ Eccentricity parameter sine component $\mathcal{U}(-1,1)$ $0.13_{-0.06}^{+0.05}$
$\sqrt{e}\cos{\omega}$ Eccentricity parameter cosine component $\mathcal{U}(-1,1)$ $0.09_{-0.45}^{+0.40}$
$e$ Eccentricity $0.09_{-0.06}^{+0.26}$
$\omega$ Argument of periastron (deg) $123_{-41}^{+105}$
$M_{p}$ Planet mass ($M_{\oplus}$) <20<20 ($2\sigma$)
$R_{p}$ Planet radius ($R_{\oplus}$) $8.2\pm 0.1$
$\rho_{p}$ Planet mean density (g/cm3) <0.2<0.2 ($2\sigma$)
$a$ Semimajor axis (au) $0.456\pm 0.005$
@@
$\ddot{\gamma}$ RV quadratic trend (m/s/day/yr) $\mathcal{U}(-1,1)$ $0.0180\pm 0.0035$
$\gamma_{\rm HARPS1}$ HARPS1 RV offset (m/s) $\mathcal{U}(25000,2700)$ $26068\pm 4$
$\sigma_{\rm HARPS1}$ HARPS1 RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $10.5_{-2.2}^{+3.0}$
$\gamma_{\rm HARPS2}$ HARPS2 RV offset (m/s) $\mathcal{U}(-100,100)$ $26095_{-10}^{+9}$
$\sigma_{\rm HARPS2}$ HARPS2 RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $8.8_{-1.8}^{+2.2}$
$\gamma_{\rm FEROS}$ FEROS RV offset (m/s) $\mathcal{U}(25000,27000)$ $26065_{-8}^{+7}$
$\sigma_{\rm FEROS}$ FEROS RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $18.7_{-5.7}^{+7.7}$
$\gamma_{\rm CORALIE}$ CORALIE RV offset (m/s) $\mathcal{U}(25000,27000)$ $26074\pm 9$
$\sigma_{\rm CORALIE}$ CORALIE RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $0.1_{-0.1}^{+2.3}$
$q_{1}^{\rm TESS}$ TESS linear limb darkening parameter $\mathcal{U}(0,1)$ $0.43_{-0.10}^{+0.11}$
$q_{2}^{\rm TESS}$ TESS quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.10_{-0.06}^{+0.08}$
$\sigma_{\rm TESS}$ TESS photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $11_{-9}^{+54}$
$q_{1}^{R}$ $R$ linear limb darkening parameter $\mathcal{U}(0,1)$ $0.53_{-0.06}^{+0.07}$
$q_{2}^{R}$ $R$ quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.92_{-0.08}^{+0.06}$
$\sigma_{\rm ASTEP}^{2023-R}$ ASTEP 2023 $R$ photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $3070_{-78}^{+75}$
$\sigma_{\rm ASTEP}^{2025-R}$ ASTEP 2025 $R$ photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $8976_{-91}^{+90}$
$q_{1}^{B}$ $B$ linear limb darkening parameter $\mathcal{U}(0,1)$ $0.23_{-0.15}^{+0.22}$
$q_{2}^{B}$ $B$ quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.15_{-0.11}^{+0.23}$
$\sigma_{\rm ASTEP}^{2023-B}$ ASTEP 2023 $B$ photometric jitter (ppm) $\mathcal{LU}(1,5\times 10^{7})$ $2257_{-99}^{+103}$
[…]
@@ 7.1 TOI-4507 b in context
TOI-4507 b appears to be an interesting planet in several ways. Figure 4 shows the results for the projected obliquity $\lambda$ along with those of other systems. The obliquity data were drawn from TEPCat [^2] https://www.astro.keele.ac.uk/jkt/tepcat/ (J. Southworth, 2011) and the age data were drawn from the NASA Exoplanet Archive (R. L. Akeson et al., 2013; J. L. Christiansen et al., 2025), both as of May 2025. Evidently, after HIP 41378 d (S. Grouffal et al., 2022), TIC 241249530 b (A. F. Gupta et al., 2024), and HD 80606 b (e.g., G. Hébrard et al., 2010). Here, we define Neptunes/sub-Saturns as the class of planets with $10<M_{p}/M_{\oplus}<90$, with the subclass of super-puffs having $\rho_{p}<0.3$ g/cm3.
TOI-4507 b joins a growing population of Neptune/sub-Saturn
Table 3: Summary of priors and posteriors of the RM effect fit. $\mathcal{U}(a,b)$ denotes a uniform prior with a start value $a$ and end value $b$. $\mathcal{G}(\mu,\sigma)$ denotes a normal prior with mean $\mu$, and standard deviation $\sigma$. $\mathcal{LU}(a,b)$ denotes a log-uniform prior with a start value $a$ and end value $b$.
Parameter Description Prior Posterior
$\lambda$ Sky-projected stellar obliquity (deg) $\mathcal{U}(-180,180)$ $15_{-44}^{+50}$
$v\sin{i_{\star}}$ Projected rotational velocity (km/s) $\mathcal{U}(0,15)$ $7.3_{-1.8}^{+2.8}$
$\rho_{\star}$ Stellar density (g/cm3) $\mathcal{G}(1.41,0.04)$ $1.41\pm 0.04$
$P$ Orbital period (days) $\mathcal{G}(104.61611,0.00006)$ $104.61611\pm 0.00006$
$t_{0}$ Transit midpoint (BJD) $\mathcal{G}(2458413.8515,0.0008)$ $2458413.8514\pm 0.0008$
$b$ Impact parameter $\mathcal{G}(0.12,0.08)$ $0.11_{-0.06}^{+0.07}$
$R_{p}/R_{\star}$ Radius ratio $\mathcal{G}(0.0723,0.0004)$ $0.0723\pm 0.0004$
$\sqrt{e}\sin{\omega}$ Eccentricity parameter sine component $\mathcal{G}(-0.13,0.06)$ $0.15\pm 0.06$
$\sqrt{e}\cos{\omega}$ Eccentricity parameter cosine component $\mathcal{G}(-0.09,0.45)$ $0.08_{-0.37}^{+0.42}$
$e$ Eccentricity $0.11_{-0.08}^{+0.20}$
$\omega$ Argument of periastrom $117_{-46}^{+96}$
$q_{1}^{\rm HARPS}$ HARPS linear limb darkening parameter $\mathcal{U}(0,1)$ $0.59_{-0.31}^{+0.27}$
$q_{2}^{\rm HARPS}$ HARPS quadratic limb darkening parameter $\mathcal{U}(0,1)$ $0.65_{-0.36}^{+0.25}$
$\beta$ Intrinsic stellar line width (km/s) $\mathcal{G}(6.0,2.0)$ $5.9_{-2.1}^{+2.0}$
$\gamma_{\rm HARPS}$ HARPS RV offset (m/s) $\mathcal{U}(25000,27000)$ $26068_{-10}^{+9}$
$\dot{\gamma}_{\rm HARPS}$ HARPS RV slope (m/s/day) $\mathcal{U}(-200,200)$ $22_{-63}^{+61}$
$\sigma_{\rm HARPS}$ HARPS RV jitter (m/s) $\mathcal{LU}(10^{-3},100)$ $1.8_{-1.8}^{+3.7}$
Additionally, with an age of $\sim 700$ Myr, TOI-4507 b is the youngest super-puff with an obliquity measurement. Further, it is one of the youngest Neptunes/sub-Saturns
@@
TOI-4507 b seems well-suited for atmospheric characterization via transit spectroscopy. Assuming zero albedo, the expected equilibrium temperature of the planet is $\sim 460$ K, lower than most other systems that have been subject to transit spectroscopy. Because of the planet’s low density, the transmission spectroscopy metric (TSM; E. M. R. Kempton et al., 2018) is $\gtrsim 160$. For reference, targets with TSM > 90 are typically considered high-quality targets for atmospheric characterization. Observations with JWST might shed light on the atmospheric composition of TOI-4507 b and the origins of its unusually low density.
@@ 7.2 Search for companions
If the RV trend is produced by a planetary companion, it should have an orbital period $P>5.6$ yr, which corresponds to the baseline of the current RV observations. The parameters of this possible companion could be constrained with additional RV and/or Gaia DR4 and DR5 astrometric measurements. When removing the signal of TOI-4507 b and the RV trend from the data, we do not see clear additional signals in the RV data, which rules out the presence of planets producing similar or larger RV signals than TOI-4507 b.